Alternative Capital Investments

The investment committee of Sentry Insurance Co. is evaluating two projects, office expansion and upgrade to computer servers. The projects have different useful lives, but each requires an investment of $1,310,000. The estimated net cash flows from each project are as follows:

Net Cash Flow
Year     

Office Expansion

    

Server

1 $366,000 $483,000
2 366,000 483,000
3 366,000 483,000
4 366,000 483,000
5 366,000
6 366,000

The committee has selected a rate of 15% for purposes of net present value analysis. It also estimates that the residual value at the end of each project's useful life is $0, but at the end of the fourth year, the office expansion's residual value would be $458,000.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162
Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Required:

If required, use the minus sign to indicate a negative net present value.

1. For each project, compute the net present value. Use the present value of an annuity of $1 table above. Ignore the unequal lives of the projects. If required, round to the nearest dollar.

Office Expansion Server Upgrade
Present value of annual net cash flows $ $
Less amount to be invested $ $
Net present value $ $

2. For each project, compute the net present value, assuming that the office expansion is adjusted to a four-year life for purposes of analysis. Use the present value of $1 table above.

Office Expansion Server Upgrade
Present value of net cash flow total $ $
Less amount to be invested $ $
Net present value $ $

3. The net present value of the two projects over equal lives indicates that the   has a higher net present value and would be a superior investment.
Answer #1
1
Office Expansion Server Upgrade
Present value of annual net cash flows 1384944 1378965
Less amount to be invested 1310000 1310000
Net present value 74944 68965
2
Office Expansion Server Upgrade
Present value of annual net cash flows 1307272 1379448
Less amount to be invested 1310000 1310000
Net present value -2728 69448
3
The net present value of the two projects over equal lives indicates that the Server Upgrade has a higher net present value and would be a superior investment.
Workings:
1
Office Expansion Server Upgrade
Present value of annual net cash flows =366000*3.784 =483000*2.855
2
Office Expansion year 4 cash flows 824000 =366000+458000
Office Expansion Server Upgrade
Year Cash flows PV factor Present value Cash flows PV factor Present value
1 366000 0.870 318420 483000 0.870 420210
2 366000 0.756 276696 483000 0.756 365148
3 366000 0.658 240828 483000 0.658 317814
4 824000 0.572 471328 483000 0.572 276276
Total 1307272 Total 1379448